OpenStudy (anonymous):

simplify no negative exponents can be used 4y–² ________ 2y–⁶

OpenStudy (anonymous):

\[\frac{2 y^6}{4 y^2}=\frac{y^4}{2} \]

OpenStudy (michele_laino):

lease keep in mind that: \[{y^{ - 2}} = \frac{1}{{{y^2}}}\]

OpenStudy (michele_laino):

and: \[{y^{ - 6}} = \frac{1}{{{y^6}}}\]

OpenStudy (anonymous):

2y^3

OpenStudy (anonymous):

\[=\frac{ 4y ^{-2} }{ 2y ^{-6} }\]

OpenStudy (anonymous):

\[=\frac{ 1 }{ 4y^2 }\div \frac{ 1 }{ 2y^6 }\]

OpenStudy (anonymous):

\[=\frac{ 1 }{ 4y^2 }\times \frac{ 2y^6 }{ 1 }\]

OpenStudy (anonymous):

\[=\frac{ 2y^6 }{ 4y^2 }\]

OpenStudy (anonymous):

simplify it @dontre

OpenStudy (anonymous):

So I would multiply to simplify? @MARC_

OpenStudy (anonymous):

Multiply it first

OpenStudy (anonymous):

if the value can be simplify,simplify it

OpenStudy (anonymous):

64/16

OpenStudy (anonymous):

@MARC_

OpenStudy (anonymous):

How u get 64/16?

OpenStudy (anonymous):

\[=\frac{ 2y^6 }{ 4y^2 }\]

OpenStudy (anonymous):

\[=\frac{ 2\times~y~\times~y~\times~y~\times~y~\times~y~\times~y }{ 4\times~y~\times~y }\]

OpenStudy (anonymous):

128/16

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

2y^4

OpenStudy (anonymous):

@robtobey ....Am I correct?